# https://ldtho.github.io/DifFUSER/

### v1

@@ Abstract
Diffusion models have recently gained prominence as powerful deep generative models, demonstrating unmatched performance across various domains. However, their potential in multi-sensor fusion remains largely unexplored. In this work, we introduce “DifFUSER”, a novel approach that leverages diffusion models for multi-modal fusion in 3D object detection and BEV map segmentation. Benefiting from the inherent denoising property of diffusion, DifFUSER is able to refine or even synthesize sensor features in case of sensor malfunction, thereby improving the quality of the fused output. In terms of architecture, our DifFUSER blocks are chained together in a hierarchical BiFPN fashion, termed cMini-BiFPN, offering an alternative architecture for latent diffusion. We further introduce a Gated Self-conditioned Modulated (GSM) latent diffusion module together with a Progressive Sensor Dropout Training (PSDT) paradigm, designed to add stronger conditioning to the diffusion process and robustness to sensor failures. Our extensive evaluations on the Nuscenes dataset reveal that DifFUSER not only achieves state-of-the-art performance with a 69.1% mIOU in BEV map segmentation tasks but also competes effectively with leading transformer-based fusion techniques in 3D object detection.
@@
Figure 1: We propose to formulate multi-modal fusion as a denoising diffusion process from noisy to refined BEV features, which are then fed into task-specific heads for 3D object detection and BEV map segmentation.
@@ 1 Introduction
3D object detection and Bird’s Eye View (BEV) map segmentation are fundamental tasks in autonomous driving. The former involves detecting objects in 3D space, while the latter segments BEV maps into semantically meaningful categories. A growing trend in this field is the fusion of features from multiple sensors, leveraging their complementary strengths. For example, 3D point clouds provide essential geometric data for 3D object detection, but lack the rich color information crucial for semantic segmentation of BEV maps—a gap effectively filled by 2D images.
@@
Recent transformer-based fusion methods [47, 1, 45] have set new benchmarks in 3D object detection by learning feature mappings from both sensors through cross-attention mechanisms. Nevertheless, these methods are less adaptable for extra tasks like BEV map segmentation. An alternative is to create unified BEV representations from both sensors, as seen in approaches like LSS [29]. However, these often yield sub-optimal results due to: 1) inadequately designed fusion module architectures that fail to capture the intricate relationship between the sensors, and 2) the inherent noise between different modalities, causing the fused features to be noisy and inaccurate.
@@
To address these challenges, we explore the use of generative models, particularly Diffusion Probabilistic Models (DPMs), for multi-modal fusion and denoising. DPMs have shown promise in various applications, including image generation [8, 35] and object detection [3]. Yet, the potential of DPMs in multi-sensor fusion, especially between 3D point clouds and 2D multi-view images, remains unexplored. To this end, we propose DifFUSER, a conditional diffusion-based generative model with enhanced fusion architecture designed for multi-modal learning in 3D perception. It processes features from both sensors to output a refined BEV representation, which is subsequently directed to task-specific heads for 3D object detection and BEV map segmentation. We also show DifFUSER’s robustness in synthesizing new features to compensate for missing modality, minimising performance loss and ensuring reliability even with compromised sensor data.
@@
In terms of architecture, our DifFUSER blocks are chained together in a hierarchical BiFPN (Bidirectional Feature Pyramid Network) fashion [37, 12], termed cMini-BiFPN, offering an alternative architecture for latent diffusion, particularly good in handling the multi-scales and highly detailed features of data from different sensors. This architectural change is complemented by a Gated Self-conditioned Modulated (GSM) latent diffusion module, designed to strengthen the conditioning modulation during diffusion, leading to more precise denoising and feature enhancement. The design and effectiveness of the DifFUSER blocks are examined in our ablation study. We train DifFUSER end-to-end using our proposed Progressive Sensor Dropout Training (PSDT) paradigm, together with task-specific losses, which enhance the model’s robustness and capability in denoising corrupted BEV features and generating more accurate outputs for downstream tasks.
@@
Fig. 2 shows the output activation maps from BEV encoder of the baseline BEVFusion [21] and our DifFUSER module, where the activation map from [21] is noisy and inaccurate, which may harm the performance of the downstream tasks. In contrast, our fusion module, with better architecture design and the denoising property of the diffusion model, can effectively mitigate the noises and intrinsic differences between different sensors, notably improving the predictions and strong robustness against sensor malfunctioning. We conduct extensive experiments on the large-scale Nuscenes dataset [2] to evaluate the performance and design choice of DifFUSER. The results show that DifFUSER achieves SOTA performance (69.1% mIOU) in the BEV map segmentation task and performs on par with SOTA transformer-based fusion methods in the 3D object detection task. We also conduct ablation studies to show the effectiveness of each component in DifFUSER.
@@
  • We are the first to introduce a new DPM-based approach for multi-modal fusion in 3D perception. With a well-designed diffusion architecture (cMini-BiFPN) and training paradigm (PSDT), our method showcases enhanced multi-modal feature fusion and denoising capabilities, as well as robustness to sensor failure by generatively synthesizing missing sensor data.
@@
  • We conduct extensive experiments on the Nuscenes dataset [2] to evaluate the performance of DifFUSER. The results show that DifFUSER surpasses the current SOTA by 6.4% in BEV map segmentation tasks and performs on par with SOTA transformer-based fusion methods in 3D object detection.
@@ 2 Related Work
In contrast, we aim to leverage the denoising property of generative models to mitigate the noises and intrinsic differences between different sensors, tackling some of the most fundamental tasks in autonomous driving, such as 3D object detection and BEV map segmentation. Our DifFUSER is pioneering in adapting the diffusion model for multi-modal fusion for 3D perception tasks. The generated BEV feature can be shared and optimized end-to-end with any potential downstream tasks.
@@ 3 Methodology
In this section, we introduce DifFUSER, a conditional generative model with enhanced fusion architecture, designed for multi-modal and multi-task learning within the scope of 3D object detection and BEV map segmentation. Our DifFUSER features takes the extracted features from lidar and camera sensors and process them in three ways: one used as diffusion target, one infused with Gaussian noise ($x_{t}$) and one is partially masked by PSDT ($\tilde{x}_{0}$). $\tilde{x}_{0}$ and $x_{t}$ are then input into the cMini-BiFPN (with GSM as encoder blocks) module, where they are denoised and refined. This enhanced representation is finally utilised by task-specific heads for precise 3D object detection and BEV map segmentation.
@@
In autonomous driving, where fast processing is the top priority, the need to ensure both the performance and efficiency of the diffusion module is important. Thus, we introduce a conditional BiFPN-like diffusion architecture, which we term as the conditional-Mini-BiFPN (cMini-BiFPN). Note that we use the term cMini-BiFPN to indicate the bi-directional feature fusion alone, without GSM module. As depicted in Fig. 3, the initial three blocks $F^{1}_{in},F^{2}_{in},F^{3}_{in}$—the encoder blocks—process both partially masked original $\tilde{x}_{0}$ and perturbed BEV features $x_{t}$, where $\tilde{x}_{0}$ serves as a guide for the diffusion process, aiding in the progressive refinement of the corrupted BEV feature. Subsequent blocks $F^{1}_{out},F^{2}_{out},F^{3}_{out}$—the decoder blocks—accept the conditioned BEV representation and apply iterative top-down and bottom-up bi-directional fusion to generate the final BEV feature. Particularly,
@@
where $\delta$ denotes the Swish activation function [37], $F^{2}_{it}$ represents the intermediate fusion result of $F^{2}_{in}$ and $F^{3}_{in}$, and $F_{out}^{2}$ is computed by blending $F^{2}_{it}$, $F_{out}^{1}$, and $F_{in}^{2}$. The weights $w_{1,2,3}^{\prime}$ and $w_{1,2,3}^{\prime\prime}$ are subject to optimisation. $F_{out}^{1}$ and $F_{out}^{3}$ are derived similarly to $F_{out}^{2}$. The final feature map is the concatenation of $F_{out}^{1}$, $F_{out}^{2}$, and $F_{out}^{3}$, which is then processed through a Transfusion-L [1] detection head and a segmentation head [21], facilitating both 3D object detection and BEV map segmentation. We believe that this streamlined and efficient cMini-BiFPN architecture is well-suited for the demanding diffusion processes in autonomous driving applications, balancing computational efficiency with the necessity for expeditious sampling.
@@ 3.2 Progressive Sensor Dropout Training (PSDT)
Sensor reliability in autonomous driving is crucial. Cameras and LiDAR systems, the eyes of robotic agents, can occasionally be occluded or fail, potentially compromising safety and operational efficiency. Recognising this, we propose the Progressive Sensor Dropout Training (PSDT) method, which enhances our DifFUSER model’s robustness and adaptability, particularly in scenarios where sensor inputs may be occluded or failed. It enables the model to generate or reconstruct missing features from camera or LiDAR inputs by leveraging the learned distribution across both modalities. This approach ensures consistent high performance, even in cases of sensor malfunction or when sensor data is partially missing, closely simulating real-world operational challenges.
Specifically in PSDT, $x_{0}$, which contains features from both camera and LiDAR, is utilised in three key ways: as a training target, as a noisy input for the diffusion module, and as a diffusion condition with randomly dropped camera or LiDAR features. To simulate the conditions of sensor dropout or malfunction, we progressively increase the dropout rate for either the camera or LiDAR inputs from 0% to a predefined maximum $\alpha=25$ during training. This is represented as:
$F_{\text{masked}}^{m}=F^{m}\odot\text{Bernoulli}(\min\left(\frac{\alpha}{100}\cdot\frac{e}{E},\frac{\alpha}{100}\right))$ (3)
@@ 3.3 Gated Self-Conditioned Modulated (GSM) Diffusion Module
Departing from traditional diffusion models [8, 35], which primarily use just $t$ as the condition, the GSM module leverages both the temporal (time step $t$) and the sensor information contained within the partially masked latent sample ($\tilde{x}_{0}$) to guide the diffusion process. This requires stronger conditioning on the noisy latent to shape the diffusion trajectory towards generating high-quality features. This dual-conditioning approach, augmented by the PSDT paradigm and task-specific losses, ensures the diffusion process focuses on feature enhancement and not mere replication, encouraging the creation of more expressive features for subsequent tasks.
Figure 4: Gated self-conditioned modulated diffusion block.
Specifically, in DifFUSER, the GSM module is used in encoder blocks $F^{1}_{in},F^{2}_{in},F^{3}_{in}$ (see Fig. 3). The GSM diffusion process is encapsulated by a FiLM-like (Feature-wise Linear Modulation) operation. This modulation computes parameters that dynamically adjust the diffusion trajectory based on the conditions $\tilde{x}_{0}$. The modulation formula is defined as:
$\tilde{x}_{0}$ $=\tilde{x}_{0}+\textit{time\_emb}(t),$ (4)
$\gamma_{t}$ $=\sigma(W_{\gamma}\cdot\tilde{x}_{0}+b_{\gamma}),$
$\alpha_{t}$ $=W_{\alpha}\cdot\tilde{x}_{0}+b_{\alpha},$
$\beta_{t}$ $=W_{\beta}\cdot\tilde{x}_{0}+b_{\beta},$
$\widetilde{x}_{t}$ $=\gamma_{t}\odot(x_{t}\odot(1+\alpha_{t})+\beta_{t}).$
Here, $\tilde{x}_{0}$ is the version of $x_{0}$ that may be partially masked due to PSDT. The parameters $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$ modulate the input feature $x_{t}$ via scaling and shifting, with $\gamma_{t}$ acting as a sigmoid-gated self-conditioning element to recalibrate layer activation based on the feature itself. These parameters are derived from the condition $x_{0}$ using a simple convolution block, we deliberately limit gradient flow when introducing noise to $x_{0}$ to prevent model memorisation. The architecture of our gated self-conditioned modulated diffusion mechanism is detailed in Fig. 4.
@@ 3.4 DifFUSER training and loss
DifFUSER is trained end-to-end in either single-task or multi-task fashion, together with task-specific losses, which encourage the module to not just denoising but also generating more expressive BEV features serving the downstream tasks. In the forward noising process, Gaussian noise $\varepsilon$ is gradually added to the extracted features $x_{0}$ via a discrete stochastic differential equation (SDE):
$q(x_{t},x_{0})=\sqrt{\bar{\alpha_{t}}}x_{0}+\varepsilon\sqrt{1-\bar{\alpha_{t}}}$ (5)
where $\bar{\alpha_{t}}$ is a hyperparameter and $\varepsilon\sim\mathcal{N}(0,I)$. If the time step $t$ is large, $x_{t}$ would become Gaussian noise.
The reverse denoising process is a discrete SDE that gradually maps a Gaussian noise into a sample. At each time step, given $x_{t}$, $t$, and $\tilde{x}_{0}$, it predicts the next reverse step:
$p^{\left(t\right)}(x_{t-1}|x_{t},\tilde{x}_{0},t)=\begin{cases}\mathcal{N}\left(p^{\left(1\right)}_{\theta}(x_{1},\tilde{x}_{0},t),\sigma^{2}I\right)&\text{if }t=1\\ q(x_{t-1}|x_{t},p^{\left(t\right)}_{\theta}(x_{t},\tilde{x}_{0},t))&\text{otherwise.}\end{cases}$ (6)
@@
Diffusion loss. Instead of using $\varepsilon$-prediction from DDPM [8], we train our model to predict the denoised and unmasked sample $x_{0}$ from the noisy $x_{t}$ at each time step, and use a simple mean-squared error loss on the prediction, thanks to the parameterisation of $\mu_{\theta}$ as a denoising network $p_{\theta}$ :
$\mathcal{L}_{diffusion}=\mathbb{E}_{t\sim[1,T],\widetilde{x}_{t}\sim q_{t}}\left[\left\|p^{\left(t\right)}_{\theta}(x_{t},t,\tilde{x}_{0})-x_{0}\right\|^{2}\right]$ (7)
where $\widetilde{x}_{0}$ is the original partially masked feature from both sensors, $x_{t}$ is the noisy feature at time step $t$, $q_{t}$ is the distribution of $x_{t}$ at time step $t$, and $p^{\left(t\right)}_{\theta}(x_{t},t,x_{0}))$ is the predicted $x_{0}$ at time step $t$.
Segmentation loss. The standard focal loss $\mathcal{L}_{seg}$ is used for the segmentation head, as in [21].
@@
Total loss. The total loss is the weighted sum of the diffusion loss, the detection loss, and the segmentation loss, depending on the training mode (single-task or multi-task):

### v2

@@ Abstract
Diffusion models have recently gained prominence as powerful deep generative models, demonstrating unmatched performance across various domains. However, their potential in multi-sensor fusion remains largely unexplored. In this work, we introduce “DifFUSER”, a novel approach that leverages diffusion models for multi-modal fusion in 3D object detection and BEV map segmentation. Benefiting from the inherent denoising property of diffusion, DifFUSER is able to refine or even synthesize sensor features in case of sensor malfunction, thereby improving the quality of the fused output. In terms of architecture, our DifFUSER blocks are chained together in a hierarchical BiFPN fashion, termed cMini-BiFPN, offering an alternative architecture for latent diffusion. We further introduce a Gated Self-conditioned Modulated (GSM) latent diffusion module together with a Progressive Sensor Dropout Training (PSDT) paradigm, designed to add stronger conditioning to the diffusion process and robustness to sensor failures. Our extensive evaluations on the Nuscenes dataset reveal that DifFUSER not only achieves state-of-the-art performance with a 70.04% mIOU in BEV map segmentation tasks but also competes effectively with leading transformer-based fusion techniques in 3D object detection.
@@
Figure 1: We propose to use denoising diffusion process for multi-modal BEV features fusion, which are then used for 3D object detection and BEV map segmentation.
@@ 1 Introduction
3D object detection and Bird’s Eye View (BEV) map segmentation are fundamental tasks in autonomous driving. The former involves detecting objects in 3D space, while the latter segments BEV maps into semantic categories. A growing trend in this field is the fusion of features from multiple sensors, leveraging their complementary strengths. For example, 3D point clouds provide geometric data for 3D object detection, but lack the rich color information crucial for BEV semantic segmentation —a gap effectively filled by 2D images.
@@
Recent transformer-based fusion methods [49, 1, 47] have set new benchmarks in 3D object detection by learning feature mappings from both sensors through cross-attention mechanisms. Nevertheless, these methods are less adaptable for extra tasks like BEV map segmentation. An alternative is to create unified BEV representations from both sensors, as seen in LSS-based [31] approaches. However, these often yield sub-optimal results due to: 1) inadequately designed fusion module architectures that fail to capture the intricate relationship between the sensors, and 2) the inherent noise between different modalities, causing the fused features to be noisy and inaccurate.
@@
To address these challenges, we explore the use of generative models, particularly Diffusion Probabilistic Models (DPMs), for multi-modal fusion and denoising. DPMs have shown promise in various applications, including image generation [9, 37] and object detection [3]. Yet, the potential of DPMs in multi-sensor fusion, especially between 3D point clouds and 2D multi-view images, remains unexplored. To this end, we propose DifFUSER, a conditional diffusion-based generative model with enhanced fusion architecture designed for multi-modal learning in 3D perception. It processes features from both sensors to output a refined BEV representation, which is subsequently directed to task-specific heads for 3D object detection and BEV map segmentation. We also show DifFUSER’s robustness in synthesizing new features to compensate for missing modality, minimising performance loss and ensuring reliability even with compromised sensor.
@@
In terms of architecture, our DifFUSER blocks are chained together in a hierarchical BiFPN (Bidirectional Feature Pyramid Network) fashion [39, 14], termed cMini-BiFPN, offering an alternative architecture for latent diffusion, particularly good in handling detailed and multi-scales features from different sensors. This architectural change is complemented by a Gated Self-conditioned Modulated (GSM) latent diffusion module, designed to strengthen the conditioning modulation during diffusion, leading to more precise denoising and feature enhancement. We train DifFUSER end-to-end using our proposed Progressive Sensor Dropout Training (PSDT) paradigm, together with task-specific losses, which enhance the model’s robustness and capability in denoising corrupted BEV features and generating more accurate outputs for downstream tasks.
@@
Fig. 2 presents the output activation maps from the BEV encoder of the baseline BEVFusion [23] and our DifFUSER module. The activation map from our BEVFusion module is blurry, which might hinder the downstream task performance. In contrast, our fusion module, with better architecture design and the denoising property of the diffusion model, can effectively mitigate the noises and intrinsic differences between different sensors, notably improving the predictions and strong robustness against sensor malfunctioning. We conduct extensive experiments on the large-scale Nuscenes dataset [2] to evaluate the performance and design choice of DifFUSER. The results show that DifFUSER achieves SOTA performance (70.04% mIOU) in the BEV map segmentation task and performs on par with SOTA transformer-based fusion methods in the 3D object detection task. We also conduct ablation studies to show the effectiveness of each component in DifFUSER.
@@
  • We are first to introduce a new DPM-based approach for multi-modal fusion in 3D perception. With a well-designed diffusion architecture (cMini-BiFPN) and training paradigm (PSDT), our method showcases enhanced multi-modal feature fusion and denoising capabilities, as well as robustness to sensor failure by generatively synthesizing missing sensor data.
@@
  • We conduct extensive experiments on the Nuscenes dataset [2] to evaluate the performance of DifFUSER. The results show that DifFUSER surpasses the current SOTA by 7.34% in BEV map segmentation tasks and performs on par with SOTA transformer-based fusion methods in 3D object detection.
@@ 2 Related Work
In contrast, we aim to leverage the denoising property of generative models to mitigate the noises and intrinsic differences between different sensors, tackling some of the most fundamental tasks in autonomous driving, such as 3D object detection and BEV map segmentation. Our DifFUSER is the first to adapt diffusion models for multi-modal fusion for 3D perception tasks. The generated BEV feature can be shared and optimized end-to-end with downstream tasks.
@@ 3 Methodology
In this section, we introduce DifFUSER, a conditional generative model with enhanced fusion architecture, designed for multi-modal and multi-task learning within the scope of 3D object detection and BEV map segmentation. Our DifFUSER features take the extracted features from lidar and camera sensors and process them in three ways: one used as the diffusion target, one infused with Gaussian noise ($x^{F}_{t}$) and one is partially masked by PSDT ($\tilde{x}^{F}_{0}$). $\tilde{x}^{F}_{0}$ and $x^{F}_{t}$ are then input into the cMini-BiFPN (with GSM as encoder blocks) module, where they are denoised and refined. This enhanced representation is finally utilised by task-specific heads for precise 3D object detection and BEV map segmentation.
@@
In autonomous driving, where fast processing is the top priority, the need to ensure both the performance and efficiency of the diffusion module is important. Thus, we introduce a conditional BiFPN-like diffusion architecture, which we term as the conditional-Mini-BiFPN (cMini-BiFPN). Note that we use the term cMini-BiFPN to indicate the bi-directional feature fusion alone, without the GSM module. As depicted in Fig. 3, the initial three blocks $F^{1}_{in},F^{2}_{in},F^{3}_{in}$—the encoder blocks—process both partially masked original $\tilde{x}^{F}_{0}$ and perturbed BEV features $x^{F}_{t}$, where $\tilde{x}^{F}_{0}$ serves as a guide for the diffusion process, aiding in the progressive refinement of the corrupted BEV feature. Subsequent blocks $F^{1}_{out},F^{2}_{out},F^{3}_{out}$—the decoder blocks—accept the conditioned BEV representation and apply iterative top-down and bottom-up bi-directional fusion to generate the final BEV feature. Particularly,
@@
where $\delta$ denotes the Swish activation function [39], $F^{2}_{it}$ represents the intermediate fusion result of $F^{2}_{in}$ and $F^{3}_{in}$, and $F_{out}^{2}$ is computed by blending $F^{2}_{it}$, $F_{out}^{1}$, and $F_{in}^{2}$. The weights $w_{1,2,3}^{\prime}$ and $w_{1,2,3}^{\prime\prime}$ are learnable and optimised during training. $F_{out}^{1}$ and $F_{out}^{3}$ are derived similarly to $F_{out}^{2}$. The final feature map is the concatenation of $F_{out}^{1}$, $F_{out}^{2}$, and $F_{out}^{3}$, which is then processed through a Transfusion-L [1] detection head and a segmentation head [23], facilitating both 3D object detection and BEV map segmentation. We believe that this streamlined and efficient cMini-BiFPN architecture is well-suited for the demanding diffusion processes in autonomous driving applications, balancing computational efficiency with the necessity for expeditious sampling.
@@ 3.2 Progressive Sensor Dropout Training (PSDT)
Sensor reliability in autonomous driving is crucial. Cameras and LiDAR systems, the eyes of robotic agents, can occasionally be occluded or fail, potentially compromising safety and operational efficiency. Recognising this, we propose the Progressive Sensor Dropout Training (PSDT) method, which enhances our DifFUSER model’s robustness and adaptability, particularly in scenarios where sensor inputs may be occluded or failed. It enables the model to generate or reconstruct missing features from camera or LiDAR inputs by leveraging the learned distribution across both modalities. This approach ensures consistently high performance, even in cases of sensor malfunction or when sensor data is partially missing, closely simulating real-world operational challenges.
Specifically in PSDT, $x^{F}_{0}$, the input for the whole fusion process, containing features from both camera and LiDAR, is utilised in three key ways: as a training target, as a noisy input for the diffusion module, and as a diffusion condition with randomly dropped camera or LiDAR features. To simulate the conditions of sensor dropout or malfunction, we progressively increase the dropout rate for either the camera or LiDAR inputs from 0% to a predefined maximum $\alpha=25$ during training. This is represented as:
$F_{\text{masked}}^{m}=F^{m}\odot\text{Bernoulli}\left(\frac{\alpha}{100}\cdot\frac{e}{E}\right)$ (3)
@@ 3.3 Gated Self-Conditioned Modulated (GSM) Diffusion Module
Departing from traditional diffusion models [9, 37], which primarily use just $t$ as the condition, the GSM module leverages both the temporal (time step $t$) and the sensor information contained within the partially masked latent sample ($\tilde{x}^{F}_{0}$) to guide the diffusion process. This requires stronger conditioning on the noisy latent to shape the diffusion trajectory towards generating high-quality features. This dual-conditioning approach, augmented by the PSDT paradigm and task-specific losses, ensures the diffusion process focuses on feature enhancement and not mere replication, encourage the synthesis of more expressive features.
Figure 4: GSM diffusion block.
Specifically, in DifFUSER, the GSM module is used in encoder blocks $F^{1}_{in},F^{2}_{in},F^{3}_{in}$ (see Fig. 3). The GSM diffusion process is encapsulated by a FiLM-like (Feature-wise Linear Modulation) operation. This modulation computes parameters that dynamically adjust the diffusion trajectory based on the conditions $\tilde{x}^{F}_{0}$. The modulation formula is defined as:
$\tilde{x}^{F}_{0}$ $=\tilde{x}^{F}_{0}+\textit{time\_emb}(t),$ (4)
$\gamma_{t}$ $=\sigma(W_{\gamma}\cdot\tilde{x}^{F}_{0}+b_{\gamma}),$
$\alpha_{t}$ $=W_{\alpha}\cdot\tilde{x}^{F}_{0}+b_{\alpha},$
$\beta_{t}$ $=W_{\beta}\cdot\tilde{x}^{F}_{0}+b_{\beta},$
$\hat{x}^{F}_{0}$ $=\gamma_{t}\cdot(x^{F}_{t}\cdot(1+\alpha_{t})+\beta_{t}).$
Here, $\tilde{x}^{F}_{0}$ is the version of $x^{F}_{0}$ that may be partially masked due to PSDT. The parameters $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$ modulate the noisy input feature $x^{F}_{t}$ via scaling and shifting, with $\gamma_{t}$ acting as a sigmoid-gated self-conditioning element to recalibrate layer activation based on the feature itself. These parameters are derived from the condition $\tilde{x}^{F}_{0}$ using a simple convolution block, we deliberately limit gradient flow when introducing noise to $x^{F}_{0}$ to prevent model memorisation. The architecture of our gated self-conditioned modulated diffusion mechanism is detailed in Fig. 4.
@@ 3.4 DifFUSER training and loss
DifFUSER is trained end-to-end in either single-task or multi-task fashion, together with task-specific losses, which encourage the module to not just denoising but also generating more expressive BEV features serving the downstream tasks. In the forward noising process, Gaussian noise $\varepsilon$ is gradually added to the extracted features $x^{F}_{0}$ via a discrete stochastic differential equation (SDE):
$q(x^{F}_{t},x^{F}_{0})=\sqrt{\bar{\alpha_{t}}}x^{F}_{0}+\varepsilon\sqrt{1-\bar{\alpha_{t}}}$ (5)
where $\bar{\alpha_{t}}$ is a hyperparameter and $\varepsilon\sim\mathcal{N}(0,I)$. If the time step $t$ is large, $x^{F}_{t}$ would become Gaussian noise (see Appendix for more discussion).
The reverse denoising process is a discrete SDE that gradually maps a Gaussian noise into a sample. At each time step, given $x^{F}_{t}$, $t$, and $\tilde{x}^{F}_{0}$, it predicts the next reverse step:
$p^{\left(t\right)}(x^{F}_{t-1}|x^{F}_{t},\tilde{x}^{F}_{0},t)=\begin{cases}\mathcal{N}\left(p^{\left(1\right)}_{\theta}(x^{F}_{1},\tilde{x}^{F}_{0},t),\sigma^{2}I\right)&\text{if }t=1\\ q(x^{F}_{t-1}|x^{F}_{t},p^{\left(t\right)}_{\theta}(x^{F}_{t},\tilde{x}^{F}_{0},t))&\text{otherwise.}\end{cases}$ (6)
@@
Diffusion loss. Instead of using $\varepsilon$-prediction from DDPM [9], we train our model to predict the denoised and unmasked sample $x^{F}_{0}$ from the noisy $x^{F}_{t}$ at each time step, and use a simple mean-squared error loss on the prediction, thanks to the parameterisation of $\mu_{\theta}$ as a denoising network $p_{\theta}$ :
$\mathcal{L}_{diffusion}=\mathbb{E}_{t\sim[1,T],\widetilde{x}^{F}_{t}\sim q_{t}}\left[\left\|p^{\left(t\right)}_{\theta}(x^{F}_{t},t,\tilde{x}^{F}_{0})-x^{F}_{0}\right\|^{2}\right]$ (7)
where $\widetilde{x}^{F}_{0}$ is the original partially masked feature from both sensors, $x^{F}_{t}$ is the noisy feature at time step $t$, $q_{t}$ is the distribution of $x^{F}_{t}$ at time step $t$, and $p^{\left(t\right)}_{\theta}(x^{F}_{t},t,x^{F}_{0}))$ is the predicted $\hat{x}^{F}_{0}$ at time step $t$.
Segmentation loss. We use focal loss $\mathcal{L}_{seg}$ for the segmentation head, as in [23].
@@
Total loss. The total loss is the weighted sum of the diffusion loss, the detection loss, and the segmentation loss, depending on the training mode: